What quantity of a 50% acid solution must be mixed with a 20% solution to produce 150mL of a 40% solution?
A. 90mL
B. 60mL
C. 50mL
D. 100mL
E. 120mL
Show work

Answer by mananth(16946)

percent ---------------- quantity
acid solution I 50.00% ---------------- x ml
acid solution II 20.00% ------ 150 - x ml
Mixture 40.00% ---------------- 150
Total 150 ml
50.00% x + 20.00% ( 150 - x ) = 40.00% * 150
50 x + 20 ( 150 - x ) = 6000
50 x + 3000 - 20 x = 6000
50 x - 20 x = 6000 - -3000
30 x = 3000
/ 30
x = 100 ml 50.00% acid solution I

m.ananth@hotmail.ca